### v2

@@ 6 Numerical illustration
$f(w)=\frac{1}{m}\sum_{i=1}^{m}\ell(y_{i}-x_{i}^{\mathsf{T}}w)+\lambda\left\|w\right\|^{2}.$ (80)
Here, $(x_{i},y_{i})\in\mathbb{R}^{n}\times\{0,1\}\ (i=1,2,\dots,m)$ denote the training example and $\ell$ is the loss function. $\lambda$ is a constant of the regularizer and is fixed to 0.01 in the numerical experiments below. We note that 80 is a type of (generalized) linear model used in the numerical experiments of [11] and [13]. As the loss function $\ell$, we use the following two functions known as robust loss functions: the Geman-McClure loss function ($\ell_{1}$) and the Cauchy loss function ($\ell_{2}$) [2] defined as
@@
Since both loss functions $\ell_{1}$ and $\ell_{2}$ are non-convex, the objective function 80 is non-convex.
@@
We test the methods on internet advertisements dataset from UCI repository[10] that is processed so that the number of instances is 600($=m$) and the number of data attributes is 1500($=n$), and the results until the stop condition $\left\|\nabla f(w_{k})\right\|<10^{-4}$ is satisfied are shown in 1. Our first observation is that RS-RNM converges faster than GD. GD does not require the calculation of Hessian or its inverse, making the time per iteration small. However, it usually needs a large number of iterations, resulting in slow convergence. Next, we look at the comparison between RNM and RS-RNM. From (a)(b) in 1, we see that RNM has the same or a larger decrease in the function value in one iteration than RS-RNM, and it takes fewer iterations to converge. This is possibly due to the fact that RNM determines the search direction in full-dimensional space. In particular, it should be mentioned that RNM converges rapidly from a certain point on, as it is shown that RNM has a super-linear rate of convergence near a local optimal solution. However, as shown in (c)(d) of 1, since RNM takes a long time to get close to the local solution along with the heavy calculation of the full regularized Hessian, RS-RNM results in faster convergence than RNM. We also confirm on figure (b) that for small dimensions $s=100,200$ a linear convergence rate seems to be achieved. However for $s=400$ it seems that the method converges super-linearly.
@@
(a) iterations versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for Geman-McClure loss
(b) iterations versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for Cauchy loss
@@
(c) computation time versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for Geman-McClure loss
@@
(d) computation time versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for Cauchy loss
Figure 1: Results by GD, RNM and RS-RNM($s=100,200,400$) under the parameter setting: $c_{1}=c_{1}^{\prime}=2,c_{2}=c_{2}^{\prime}=1,\gamma=\gamma^{\prime}=0.5,\alpha=0.3,\beta=0.5$ for RNM and RS-RNM.
@@ 7 Conclusions
Random projections have been applied to solve optimization problems in suitable lower-dimensional spaces in various existing works. In this paper, we proposed the randomized subspace regularized Newton method (RS-RNM) for a non-convex twice differentiable function in the expectation that a framework for the full-space version [26, 27] could be used; indeed, we could prove the stochastic variant of the same order of iteration complexity, i.e., the global complexity bound of the algorithm: the worst-case iteration number $m$ that achieves $\min_{k=0,\ldots,m-1}\left\|\nabla f(x_{k})\right\|\leq\varepsilon$ is $O(\varepsilon^{-2})$ when the objective function has Lipschitz Hessian. On the other hand, although RS-RNM uses second-order information similar to the regularized Newton method having a super-linear convergence, we proved that it is not possible, in general, to achieve local super-linear convergence and that local linear convergence is the best rate we can hope for in general. We were however able to prove super-linear convergence in the particular case where the Hessian is rank deficient at a local minimizer. In this paper we choose to thoroughly investigate local convergence rate for the Newton-based method. One could possibly, in a future work, extend these results to a state-of-the-art second order iterative method and compare the resulting subspace method with other state-of-the-art algorithms.

### v3

@@ 6.1 Support vector regression
(92) $f(w)=\frac{1}{m}\sum_{i=1}^{m}\ell(y_{i}-x_{i}^{\mathsf{T}}w)+\lambda\left\|w\right\|^{2}.$
Here, $(x_{i},y_{i})\in\mathbb{R}^{n}\times\{0,1\}\ (i=1,2,\dots,m)$ denote the training example and $\ell$ is the loss function. $\lambda$ is a constant of the regularizer and is fixed to 0.01 in the numerical experiments below. We note that 92 is a type of (generalized) linear model used in the numerical experiments of [18] and [22]. As the loss function $\ell$, we use the following two functions known as robust loss functions: the Geman-McClure loss function ($\ell_{1}$) and the Cauchy loss function ($\ell_{2}$) [2] defined as
@@
Since both loss functions $\ell_{1}$ and $\ell_{2}$ are non-convex, the objective function 92 is non-convex.
@@
We test the methods on internet advertisements dataset from UCI repository [15] that is processed so that the number of instances is 600(=m=m) and the number of data attributes is 1500(=n=n), and the results, until the stop condition $\left\|\nabla f(w_{k})\right\|<10^{-4}$ is satisfied, are shown in Figures 2, 4, 2, and 4. Our first observation is that RS-RNM converges faster than GD. GD does not require the calculation of Hessian or its inverse, making the time per iteration small. However, it usually needs a large number of iterations, resulting in slow convergence. Next, we look at the comparison between RNM and RS-RNM. From Figures 2 and 4, we see that RNM has the same or a larger decrease in the function value in one iteration than RS-RNM, and it takes fewer iterations to converge. This is possibly due to the fact that RNM determines the search direction in full-dimensional space. In particular, it should be mentioned that RNM converges rapidly from a certain point on, as it is shown that RNM has a super-linear rate of convergence near a local optimal solution. However, as shown in Figures 2 and 4, since RNM takes a long time to get close to the local solution due to the heavy calculation of the full regularized Hessian, RS-RNM results in faster convergence than RNM. We also confirm on Figure 4 that for small dimensions $s=100,200$ a linear convergence rate seems to be achieved. However for $s=400$ it seems that the method converges super-linearly.
@@
Figure 1: Iterations versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for Geman-McClure loss
@@
Figure 2: Computation time versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for Geman-McClure loss
@@
Figure 3: Iterations versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for Cauchy loss
Figure 4: Computation time versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for Cauchy loss
6.2 Low rank Rosenbrock function
To properly illustrate the superlinear convergence proved in the low rank setting (c.f. Section 5.3), we conducted numerical experiments on a low rank Rosenbrock function: $f(x)=R(U^{\top}Ux)$, where
$R(x)=\sum_{i=1}^{n-1}100(x_{i+1}-x_{i}^{2})^{2}+(x_{i}-1)^{2},$
and $U\in\mathbb{R}^{r\times n}$ is a matrix whose columns are orthogonal. If we denote by $\Pi\in\mathbb{R}^{n\times n}$ the matrix $U^{\top}U$, we see that for all $x\in\mathbb{R}^{n}$, $f(x)=f(\Pi x)$, hence the Hessian of $f$ is of rank $r$ for all $x\in\mathbb{R}^{n}$. The parameters in Section 3 are fixed as follows:
$c_{1}=c_{1}^{\prime}=2,c_{2}=c_{2}^{\prime}=1,\gamma=\gamma^{\prime}=0.5,\alpha=0.3,\beta=0.5,s\in\{100,200,600\}.$
Figure 5: Iterations versus $\left\|\nabla f(x)\right\|$ ($\log_{10}$-scale) for low rank Rosenbrock function
Figure 6: Computation time versus $\left\|\nabla f(x)\right\|$ ($\log_{10}$-scale) for low rank Rosenbrock function
Figures 6 and 6 show experiments for $n=3000$ and $r=500$. We selected three values for $s$, two ($s=100,200$) smaller than $r$ and one ($s=600$), larger than $r$. The results confirm the results of Section 5: when $s>r$ we have local superlinear convergence, otherwise the convergence is only linear locally.
6.3 Convolutional neural network
We tested our method on a micro Convolutional Neural Network (CNN) using the MNIST dataset in [13]. We used the cross-entropy loss function $m=256$ images. Our CNN is made of the following factors:
  • one convolutional layer (1 input channel, 1 output channel, kernel size 3),
  • a ReLU activation,
  • a max pooling layer (kernel size 2),
  • a fully connected layer mapping the flattened feature vector to 10 classes.
This setup is intended to demonstrate the differences between the three methods in a controlled, small-scale scenario. This problem is formulated as
$\min_{w\in\mathbb{R}^{n}}\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\mathcal{M}(w,x_{i}),y_{i}),$
where $(x_{i},y_{i})$ denotes the MNIST dataset with $x_{i}\in\mathbb{R}^{784}$ and $y_{i}\in\{0,1\}^{10}$ ($m=256$), $\mathcal{L}$ denotes the Cross Entropy Loss function, and $\mathcal{M}$ denotes the CNN with $n=1710$ parameters. The parameters in Section 3 are fixed as follows:
$c_{1}=c_{1}^{\prime}=2,c_{2}=c_{2}^{\prime}=1,\gamma=\gamma^{\prime}=0.5,\alpha=0.3,\beta=0.5,s\in\{100,200,500\}.$
The results are show in Figures 8 and 8. We notice that our method outperforms GD which is stuck at some stationary point and RNM which is to slow to converge.
Figure 7: Iterations versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for CNN with the MNIST dataset
Figure 8: Computation time versus $\left\|\nabla f(w)\right\|$ ($\log_{10}$-scale) for CNN with the MNIST dataset
6.4 Choice of $s$
In the special case where the Hessian truly has low-rank structure, setting $s$ to this value can substantially speed up convergence, provided the rank is not prohibitively large. However, in more general problems, especially where the Hessian does not exhibit pronounced low-rank properties or its effective rank is unknown, preselecting $s$ is more challenging. One might try to start with some constant value of $s$ and increasing it gradually since the best $s$ ultimately depends on problem-specific characteristics and computational resources.
@@ 7 Conclusions
Random projections have been applied to solve optimization problems in suitable lower-dimensional spaces in various existing works. In this paper, we proposed the randomized subspace regularized Newton method (RS-RNM) for a non-convex twice differentiable function in the expectation that a framework for the full-space version [39, 40] could be used; indeed, we could prove the stochastic variant of the same order of iteration complexity, i.e., the global complexity bound of the algorithm: the worst-case iteration number $m$ that achieves $\min_{k=0,\ldots,m-1}\left\|\nabla f(x_{k})\right\|\leq\varepsilon$ is $O(\varepsilon^{-2})$ when the objective function has Lipschitz Hessian. On the other hand, although RS-RNM uses second-order information similar to the regularized Newton method having a super-linear convergence, we proved that it is not possible, in general, to achieve local super-linear convergence and that local linear convergence is the best rate we can hope for in general. We were however able to prove super-linear convergence in the particular case where the Hessian is rank deficient at a local minimizer. In this paper we choose to thoroughly investigate local convergence rate for the Newton-based method. One could possibly, in a future work, extend these results to a state-of-the-art second order iterative method and compare the resulting subspace method with other state-of-the-art algorithms, as [19, 47, 48].